Samantha is 4 years older than her brother. The sum of their ages is 20. The product of their ages is 96. How old is Samantha's brother?
Answer by JulietG(1812)

S = B+4
S + B = 20
S * B = 96
Substitute the known value of S in the first equation into the second.
(B+4) + B = 20
Add
2B + 4 = 20
Subtract 4 from each side
2B = 16
Divide each side by 2
B = 8
If B is 8, then S is 8+4, or 12.
8*12=96
Success!